OpenStudy (anonymous):

hey can anyone help me find the y and x intercepts for f(x)=x^2+3x+40

OpenStudy (anonymous):

the x intercept is where f(x) equals zero so 0=x^2+3x+40 if you use the x= -b+/- \[\sqrt{b^2 - 4ac} and divide all that by 2a you get your answer. for y intercept is where x equals zero so f(x) = (0)^2 +3(0)+ 40 so f(x) = 0 + 0+ 40 f(x) = 40

OpenStudy (anonymous):

I dont understand what you just wrote.....the question ask y=__________x=___________

OpenStudy (anonymous):

Type an ordered pair

OpenStudy (anonymous):

i don't want to give you the answer, it won't help you learn. do you follow this? f(x) can be considered as y. so if x=0 for your y-intercept. then you will get y = 0^2+ 3(0) + 40 which ends up giving you y= 40

OpenStudy (anonymous):

for x-intercept, y must equal 0 using quadratic formula you get... x= -3 \[\pm\] \[\sqrt{3^2 -4(1)(40)}\] and then divide all of it by 2

OpenStudy (anonymous):

are you sure thats your equation, not -40 ?